# Growing Trees and Amoebas’ Replications

### v1

@@ Abstract
An amoeba is a tree with a number assigned to each vertex. We describe a natural process of growing trees from a given amoeba and discuss conditions for such a process to be finite.
@@ 1 Introduction
We describe a process of growing trees from an amoeba, which is a tree $H$ with a number assigned to each vertex. In a simplified language, this number $m(v)$ tells how many extra edges we will require around a $v\in V(H)$. Given a tree $T$ we look at a fixed copy $H^{\prime}$ of $H$ within $T$, and look whether every $v\in V(H^{\prime})$ is incident with at least $m(v)$ edges within $E(T)\setminus E(H^{\prime})$. If not, we attach the corresponding number of pendent edges to $v$, increasing the size of $T$. A particularly interesting question is whether such growth process can stop or go on forever. This has relations to several known models:
String and graph generation models
@@
are intensively studied in theoretical computer science, formal language theory, and mathematical biology. There are many models of this sort: cellular automata, rewriting systems, graph grammars etc. [13, 6, 10, 12]. In particular, our model resembles graph relabelling systems with forbidden contexts [12]. Most of graph generation models use labelling of graphs. In contrast, we propose a simple rule of tree growth that governed by the structure of subtrees and does not use labels on vertices and/or edges. The resulting process is interesting from graph-theoretical point of view and possibly might impact the area of graph generation models.
Galton-Watson process
@@
is a foundational concept in the realm of probability theory, specifically in the study of family trees and the dynamics of progeny within populations [7]. Introduced by Francis Galton and Henry Watson [3, 7], this stochastic process serves as a mathematical model to examine the transmission of traits, characteristics, or lineage through successive generations of families [7, 11]. The key idea revolves around the probability distribution of offspring within a family, where individuals independently give rise to a random number of descendants [7]. This model has proven particularly valuable in fields such as genetics and demography [3], providing insights into the likelihood of lineage continuation or extinction [7]. The Galton-Watson process remains an essential tool for understanding the unpredictable nature of family structures and the variability in the number of progeny across generations [11].
@@
In contrast to the probabilistic nature of the Galton-Watson process, our process represents a deterministic analog with a distinct feature—its replication criteria are not dictated by a random probability distribution. Instead, the expansion is orchestrated by the existence of a precisely chosen subgraph within the existing tree, acting as a deterministic trigger for the generation of additional leaves or progeny. This deterministic analog introduces a structured and controlled element to the reproductive dynamics, offering a unique perspective on how specific configurations within a family lineage can influence and determine the proliferation of descendants. This departure from stochastic randomness to a deterministic criterion opens up new avenues for exploring the interplay between structure and growth of trees, providing a nuanced understanding of the factors that contribute to the perpetuation or divergence of lineages.
Avoidable paths
were studied in [9, 2, 1]. This concept was generalized further to the case of arbitrary two-rooted graphs in [8] where the so-called inherent graphs were introduced. Verifying inherence is a difficult task. For example, it is not known whether the path $P_{3}$ with edges $st$ and $tv$ is inherent provided $s$, $t$ are the roots. Interestingly, this problem is related to the cops and robbers game [5]. The paper [8] presents necessary conditions for inherence in terms of a pendant extension process, which extends a two-rooted graph by pendant edges.
The process of growing trees from an amoeba generalizes the pendant extension process. It might be useful for further generalizations of avoidability.
@@
This paper is structured as follows. In the beginning of Section 2 we introduce the main concepts of the paper, such as amoeba, its extension, growth, mortality and confinement. In Section 2.1 we relate mortality and confinement conditions while in Section 2.2 we discuss relations between the symmetries of the underlying graphs and the amoeba mortality. In Section 3 we focus on the mortality conditions for amoebas growing in a simplest way, namely, by attaching pendant edges. We identify degree conditions which are necessary for any such amoeba to be immortal in Section 3.1, while in Section 3.2 we describe the immortal amoebas whose underlying graph is a caterpillar. We conclude the paper by Section 4, in which we propose some conjectures as well as directions for future work.
@@ 2 Amoebas
An ordinary amoeba [^1] This name was recently given to a different class of graphs [4]. We think that these two classes can not be confused. $A=(T,m)$ is defined as a pair consisting of a tree $T$, together with the root function $m:V(T)\to\mathbb{N}$ mapping every vertex $v$ of $T$ to a non-negative integer $m(v)$ called multiplicity of $v$.
A vertex from $A$ of positive multiplicity is called a root of $A$. Furthermore define $k(A)=\sum_{v\in V(H)}m(v)$.
If not stated otherwise, the amoebas discussed in this paper will always be ordinary.
@@
A copy of amoeba $(H,m)$ in a graph $G$ is an amoeba $(H^{\prime},m^{\prime})$ where $H^{\prime}\leq G$, and $H\cong H^{\prime}$ such that the functions $m,m^{\prime}$ coincide under the corresponding isomorphism. When no fear of ambiguity, we denote amoeba and its copy by the same notation.
Let $A$ be a copy of an amoeba in a tree $G$. Then a $\ell$-growth $G^{\prime}$ of $A$ in $G$ is a tree containing both $G$ and an $\ell$-extension of $A$ and having the minimal possible number of edges.
@@
Note that a copy of an amoeba may have non-isomorphic growths, see example in Fig. 1.
Figure 1: Non-isomorhic 3-growths of the same copy of amoeba. The copy is shadowed in grey, the new edge is dashed
For an amoeba $A=(H,m)$, a $\ell$-growth sequence of $A$ starting at a tree $T$ is a sequence of trees $T_{0},T_{1},\dots$ such that $T_{0}=T$ and $T_{i+1}$ is a growth of a copy of $A$ in $T_{i}$, and $T_{i+1}\neq T_{i}$.
For an amoeba $A=(H,m)$, a $\ell$-generation sequence of $A$ is a maximal $\ell$-growth sequence starting at $H$.
If $\ell$ is clear from the context it will be omitted in notation for extensions, growths, generation sequences etc.
Let $\ell$ be the extension parameter and $A$ be an amoeba. It is immortal if it admits an infinite generation sequence, otherwise it is mortal. Amoeba $A$ is strongly immortal if all its generation sequences are infinite. Amoeba $A$ is partially immortal if it admits an infinite growth sequence.
A tree $T$ is confining for an amoeba $A=(H,m)$ if it admits a copy of $A$ and all copies of $A$ in $T$ are dead.
@@
In [8], general two-rooted graphs are considered in the extension process. Since attaching paths to a graph does not change its set of cycles, this process is finite for non-trees. As we mainly study mortality, we restrict our attention to the case of trees.
@@ 2.1 Initial observations
Note that a copy of an amoeba may have non-isomorphic growths, which make things complicated. So, we start from study of this effect to show that it appears for small amoebas (with respect to $\ell$) and in early members of growth sequences. An example in Fig. 1 can not be seen in generation sequences, as we will prove. It is caused by the minimality condition.
Let $T^{\prime}$ be an $\ell$-growth of a alive copy of an amoeba $A=(H,m)$ in a tree $T$, and let $A^{\prime}$ be the corresponding extension of $A$ contained in $T^{\prime}$, and $P$ be a path $v_{0},v_{1},\dots,v_{\ell}$ of the extension starting at a vertex $v_{0}$ such that $\{v_{\ell-1},v_{\ell}\}\notin E(G)$. Take a connected component of the forest $T^{\prime}-H$ containing $v_{1}$. It is a tree $T^{\prime\prime}$ rooted at $v_{1}$, which is called an area of $P$. Due to the minimality condition, the depth of $T^{\prime\prime}$ (the maximal length of a path from the root to a leaf) is $\ell-1$. Moreover, $v_{1},\dots,v_{\ell}$ is a unique path of length $\ell-1$ from the root $v_{1}$. Note that other edges in $T^{\prime}$ are the edges of the original tree $T$. In this way we come to the following conlusion.
@@
An intersection of an area of an extension path and the original graph is a tree of depth at most $\ell-1$.
Given a tree $T$ and its subtree $T^{\prime}$, there is a partial order on $V(T)\setminus V(T^{\prime})$ defined as follows: $u\leq^{T^{\prime}}v$ if $u$ belongs to the path from $v$ to $T^{\prime}$. For each $v\in V(T)\setminus V(T^{\prime})$ the set $\{u:v\leq^{T^{\prime}}u\}$ induces a tree ${\mathcal{I}}_{v}^{T^{\prime}}$ rooted at $v$. We omit the mentioning of $T^{\prime}$ in the exponent when it is clear from context.
@@
Let $T_{0},T_{1},\dots,T_{k}$ be an $\ell$-growth sequence for an amoeba $A=(H,m)$ and $v\in V(T_{k})\setminus V(T_{0})$ such that ${\mathcal{I}}_{v}^{T_{0}}$ has the depth at most $\ell-1$. If for some vertex $u$ in ${\mathcal{I}}_{v}^{T_{0}}$ there are two different paths $P$ and $P^{\prime}$ of equal lengths from $u$ to a leaf, then
[…]
@@
W.l.o.g. we can assume that $P$ and $P^{\prime}$ have no common edges. Indeed, otherwise we can change the vertex $u$ by the last common vertex of $P$ and $P^{\prime}$. Choose the maximal $i$ such that at least one edge of $P\cup P^{\prime}$ is not in $T_{i}$. W.l.o.g. we can assume that $e=(u_{0},u_{1})$ is the last edge of $P$ and $e$ is not in $T_{i}$. It means that $e$ is added in a growth in $T_{i}$ of a copy $A=(H,m)$ of the amoeba. So an extension of $A$ contains an extension path from a root $r$ of $A$ to $u_{1}$. We claim that $u$ is not on the path from $r$ to $u_{1}$ that has no common edges with $P^{\prime}$. Indeed, otherwise $P^{\prime}$ does not intersect $H$ and the extension path can be changed to the path containing $P^{\prime}$, contradicting the minimality condition, see Fig. 2(a). Also, $r$ is not on $P$. Indeed, otherwise the extension should add more edges, since the length of the whole path $P$ is less than $\ell$, see Fig. 2(b).
@@
  • vertices $u$, $u_{1}$, $r$ are end-points of paths starting at a vertex $w$, Fig. 2(c);
  • the path from $r$ to $u_{1}$ has a common edge with $P^{\prime}$, see Fig. 2(d).
In both cases $A$ is contained in the subtree rooted at $r$ since the path from $r$ to $u_{1}$ has no common edges with $A$. This subtree is inside ${\mathcal{I}}^{T_{0}}_{u}$ so its depth is at most $\ell-2$, which implies that the diameter of $H$ is at most $42\ell-$ 4. Also, any path starting at $r$ and containing $w$ and $u$ should has the length at most $\ell-1$, by the minimality condition. Since $T_{i}$ contains $T_{0}$ and the path from $r$ to $T_{0}$ has at least 3 edges, the diameter of $T_{0}$ is at most $62\ell-$ 6. ∎
Lemma 2 puts restrictions on extensions using ‘new’ edges only. Ambiguous $\ell$-growths are possible in generation sequences even if $H$ has the diameter at least $62\ell-$ 6. See example in Fig. 3.
@@
It is open if these three versions of immortality are equivalent.
Figure 4: An immortal amoeba with one root with respect to $2$-extensions. Next step uses the left lower $P_{5}$ for extensions. Repetitions of the process continue forever.
It is clear that the last tree in any generation sequence is confining. So mortality implies the existence of confining tree. We do not know whether the opposite is true. We are able to prove the opposite for the values of the extension parameter $\ell\in\{1,2\}$.
@@
In one direction it is obvious.
In other direction: take a confining tree and a copy of the amoeba. Start a generation sequence from this copy. Note that $1$- and $2$-growths of a copy of the amoeba in any tree are isomorphic since in a rooted tree of depth $\leq 1$ all leaves belong to the same orbit of automorphisms of the tree. It implies that at each step of extension new edges can be chosen from the confining tree. Therefore the sequence is finite. ∎
@@
In general case, strong immortality implies immortality but the converse is open. For the extension parameterr $\ell\in\{1,2\}$ the converse follows from Lemma 3. If an amoeba is not strongly immortal then the last tree of its finite generation sequence is a confining tree for the amoeba. Thus any its generation sequence is finite.
In the proof of the next theorem we will use the notion of full $t$-ary tree of depth $\ell$. To this end we denote by $T^{t,\ell}$ a tree where every node on depth less than $\ell$ has $t$ children while every node on depth $\ell$ is a leaf.
@@
An amoeba is not partially immortal if and only if it is totally confined, regardless of its extension parameter.
@@
In one direction it is easy. Suppose an amoeba is not partially immortal. Take a tree $T$ and any $\ell$-growth sequence $T=T_{0},\dots,T_{k}$ starting at $T$. It is finite and the last tree $T_{k}$ in the sequence confines the amoeba and contains $T$.
For the other direction, take a tree $T_{0}$ and a totally confined amoeba $A=(H,m)$. Let $\delta=\max_{v\in V(H)}\big{(}d(v)+m(v)\big{)}$, $d=\mathop{\mathrm{diam}}(T_{0})$. Choose a confining tree $T$ that contains a full $\delta$-ary tree $T^{\delta,3\ell+2d}$ of sufficiently large depth, say, $3\ell+2d$. Choose a copy of $T_{0}$ containing the center of $T^{\delta,3\ell+2d}$. Start $\ell$-growth sequence $T_{0},T_{1},\dots$ from this particular copy. If $\mathop{\mathrm{diam}}(T_{i})\leq 2\ell-6$ then any copy of $A$ in it is at distance at most $62\ell-$ 6 from the center of $T^{\delta,3\ell+2d}$. Thus any growth of $T_{i}$ can be embedded to $T$. If an $\ell$-growth of $T_{i}$ uses edges of $T_{0}$, then the corresponding copy of $A$ is at distance at most $2d$ from the center of $T^{\delta,3\ell+2d}$. Thus, again, the growth can be embedded to $T$. If an $\ell$-growth of $T_{i}$ does not use edges of $T_{0}$ and $\mathop{\mathrm{diam}}(T_{i})>2\ell-6$, then Lemma 2 implies that all such growths are isomorphic. So we apply the same arguments as in the proof of Lemma 3 and conclude that $T_{i+1}$ can be embedded in $T$. Therefore the sequence is finite. ∎
Conjecture 1.
Let $\ell$ be the extension parameter, let $A=(H,m)$ be an amoeba. If $\Delta(H)>k(A)+2$ then $A$ is totally confined.

### v2

@@ Abstract
An amoeba is a tree together with instructions how to iteratively grow trees by adding paths of a fixed length $\ell$. This paper analyses such a growth process. An amoeba is mortal if all versions of the process are finite, and it is immortal if they are all infinite.
We obtain some necessary and some sufficient conditions for mortality. In particular, for growing caterpillars in the case $\ell=1$ we characterize mortal amoebas. We discuss variations of the mortality concept, conjecture that some of them are equivalent, and support this conjecture for $\ell\in\{1,2\}$.
Keywords: tree, amoeba, caterpillar, mortality
@@ 1 Introduction
A classical result of Dirac from 1961 [10] states that every (non-null) chordal graph has a vertex whose neighborhood is a clique. This result was generalized in the literature in various ways [16, 2, 3, 8, 4, 17]. In 2019, Beisegel et al. [1] proposed a common generalization of all these results by introducing the concept of avoidable paths. An induced path $P$ in a graph $G$ is said to be avoidable if every extension of $P$ is contained in an induced cycle. Here, an extension of $P$ is any induced path in $G$ that can be obtained by extending $P$ by one edge from each endpoint. Beisegel et al. conjectured that for every positive integer $k$, every graph that contains an induced $k$-vertex path also contains an avoidable induced $k$-vertex path, and proved the statement for the case $k=2$. The general conjecture was proved in 2020 by Bonamy et al. [5]. A further strengthening was given by Gurvich et al. [12], who showed that in every graph, every induced path can be transformed into an avoidable one via a sequence of shifts (where two induced $k$-vertex paths are said to be shifts of each other if their union is an induced path with $k+1$ vertices).
In [13] the concept of avoidability was applied to general graphs rather than only paths. A graph $H$ with roots $s,t\in V(H)$ is said to be inherent if every graph $G$ that contains an induced copy of $H$, also contains an avoidable induced copy of $H$. In the special case when $H=(v_{1},\dots,v_{n})$ is a path with $s=v_{1}$ and $t=v_{n}$, the result of Bonamy et al. [5] states that $(H,s,t)$ is inherent. In [13] the authors develop a so-called pendant extension method for showing that a fixed $2$-rooted graph is not inherent, and identify many such $2$-rooted graphs. This method is based on the process of growing trees. [^1] The case of trees is the only non-trivial case: using the pendant extension method it is easily shown that the process stops for all connected graphs which contain a cycle (see also Remark 2). If this process halts, a particular $2$-rooted tree is non-inherent. It is open if the halting problem for this process is decidable.
The above process is of independent interest. In this paper we generalize it in two ways. First, we allow arbitrarily many roots. Second, a step of our process may add an arbitrary path to a root, rather than a single edge.
An amoeba is a tree with a number of roots assigned to each vertex. A tree $T$ may have several copies of a particular amoeba. Each one requires a ‘certain growth space’ in $T$. If it cannot grow within $T$, the amoeba will live on by expanding $T$ – attaching new paths at the points of its roots. In this way new copies of the amoeba may appear in the expanded $T$. Alternatively, if every copy of the amoeba in $T$ has enough growth space, the $T$ will not grow anymore. We say that $T$ confines such an amoeba. To specify the finiteness of growth of a given amoeba we introduce several notions of mortality. In this paper we present some necessary and some sufficient conditions for amoeba mortality.
Our process of growing trees has relations to several known models:
@@
String and graph generation models    are intensively studied in theoretical computer science, formal language theory, and mathematical biology. There are many models of this sort: cellular automata, rewriting systems, graph grammars, etc. [19, 9, 14, 18]. In particular, our model resembles graph relabelling systems with forbidden contexts [18]. Most graph generation models use labelling of graphs. In contrast, we propose a simple rule of tree growth governed by the structure of subtrees and does not use labels on vertices and/or edges. The resulting process is interesting from the graph-theoretical point of view and possibly might impact the area of graph generation models.
@@
The Galton-Watson process    is a probabilistic process of growing trees. It is a foundational concept in the realm of probability theory, specifically in the study of family trees and the dynamics of progeny within populations [11]. Introduced by Francis Galton and Henry Watson [6, 11], this stochastic process serves as a mathematical model to examine the transmission of traits, characteristics, or lineage through successive generations of families [11, 15]. The key idea revolves around the probability distribution of offspring within a family, where individuals independently give rise to a random number of descendants [11]. This model has proven particularly valuable in fields such as genetics and demography [6], providing insights into the likelihood of lineage continuation or extinction [11]. The Galton-Watson process remains an essential tool for understanding the unpredictable nature of family structures and the variability in the number of progeny across generations [15].
@@
In contrast to the Galton-Watson process, the replication criteria of the amoeba replication process is not dictated by a probability distribution. Instead, the expansion is orchestrated by the existence of a specific subgraph within the existing tree, acting as a trigger for the generation of additional leaves or progeny. This departure from stochastic randomness to a combinatorial criterion opens up new avenues for exploring the interplay between structure and growth of trees.
@@
This paper is structured as follows. In the beginning of Section 2 we introduce the main concepts of the paper, such as amoeba, its extension, growth, mortality, immortality, and confinement. In Section 2.1 we relate mortality and confinement conditions, while in Section 2.2 we discuss relations between the symmetries of the underlying graphs and amoeba mortality. In Section 3 we focus on the mortality conditions for amoebas growing in the simplest way, namely, by attaching pendant edges. We identify degree conditions which are necessary for any such amoeba to be not mortal in Section 3.1, while in Section 3.2 we describe the not mortal amoebas whose underlying graph is a caterpillar. We conclude the paper with Section 4, in which we propose some conjectures as well as directions for future work.
@@ 2 Amoebas
An amoeba [^2] This name was recently given to a different class of graphs [7]. We think that these two classes cannot be confused. $A=(T,m)$ is defined as a pair consisting of a tree $T$, together with the root function $m:V(T)\to\mathbb{N}$ mapping every vertex $v$ of $T$ to a non-negative integer $m(v)$ called the multiplicity of $v$.
A vertex from $A$ of positive multiplicity is called a root of $A$. Furthermore, we define the number of roots of $A$ to be $\sum_{v\in V(H)}m(v)$. The set of roots of $A$ is denoted by $\operatorname{supp}(m)$.
@@
A copy of amoeba $(H,m)$ in a graph $G$ is an amoeba $(H^{\prime},m^{\prime})$ where $H^{\prime}\leq G$, and $H\cong H^{\prime}$ such that the functions $m,m^{\prime}$ coincide under the corresponding isomorphism. When there is no possibility of ambiguity, we denote an amoeba and its copy by the same notation.
Let $A$ be a copy of an amoeba in a tree $G$. Then an $\ell$-growth $G^{\prime}$ of $A$ in $G$ is a tree containing both $G$ and an $\ell$-extension of $A$ and having the minimal possible number of edges. Note that an $\ell$-growth is not uniquely defined.
@@
Note that a copy of an amoeba may have non-isomorphic growths; see an example in Fig. 1.
Figure 1: Non-isomorhic 3-growths of the same copy of an amoeba. The copy is shadowed in grey, the new edge is dashed
For an amoeba $A=(H,m)$, an $\ell$-growth sequence of $A$ starting in a tree $T$ is a (possibly infinite) sequence of trees $T_{0},T_{1},\dots$ such that $T_{0}=T$ and $T_{i+1}$ is an $\ell$-growth of a copy of $A$ in $T_{i}$, and $T_{i+1}\neq T_{i}$. An $\ell$-growth sequence is maximal if it cannot be extended.
It will be useful to define the following special case of maximal $\ell$-growth sequences: for an amoeba $A=(H,m)$, an $\ell$-generation sequence of $A$ is a maximal $\ell$-growth sequence starting in $H$. See an example in Fig. 2.
Figure 2: A generation sequence for an amoeba $H$ with $\ell=2$. The $\ell$-growth of $T_{4}$ is $T_{4}$
If $\ell$ is clear from the context it will be omitted in notation for extensions, growths, generation sequences, etc.
Let $\ell$ be the extension parameter, let $A$ be an amoeba, and let $T$ be a tree. We introduce the following definitions.
  • Amoeba $A$ is immortal w.r.t $T$ if all maximal $\ell$-growth sequences of $A$ starting in $T$ are infinite.
  • Amoeba $A$ is mortal w.r.t $T$ if all maximal $\ell$-growth sequences of $A$ starting in $T$ are finite.
  • When $A=(H,m)$ is (im)mortal w.r.t $H$, we say that $A$ is (im)mortal.
  • A tree $T$ is confining for an amoeba $A$ if it admits a copy of $A$ and all copies of $A$ in $T$ are dead.
Note that any $\ell$-growth sequence starting in a confining tree is of length $1$. It remains open whether all generation sequences of $A$ can be infinite.
@@
One can use general graphs rather than trees in the extension process (in [13] it was done for general two-rooted graphs). Since attaching paths to a graph does not change its set of cycles, this process is finite for non-trees. As we mainly study mortality, we restrict our attention to the case of trees.
@@ 2.1 Initial observations
Note that a copy of an amoeba may have non-isomorphic growths, which lead to essentially different growth sequences. We are going to show that this effect can take place only for early stages of the growth process and only for amoebas that are small enough w.r.t the growth parameter $\ell$. An example in Fig. 1 cannot be seen in generation sequences, as we will prove. It is caused by the minimality condition.
Let $T^{\prime}$ be an $\ell$-growth of an alive copy of an amoeba $A=(H,m)$ in a tree $T$, and let $A^{\prime}$ be the corresponding extension of $A$ contained in $T^{\prime}$, and $P$ be a path $v_{0},v_{1},\dots,v_{\ell}$ in the extension starting in a vertex $v_{0}\in V(H)$ such that $\{v_{\ell-1},v_{\ell}\}\notin E(T)$. Take a connected component of the forest $T^{\prime}-H$ containing $v_{1}$. It is a tree $T^{\prime\prime}$ rooted at $v_{1}$, which is called an area of $P$. Due to the minimality condition, the depth of $T^{\prime\prime}$ (the maximal length of a path from the root to a leaf) is $\ell-1$. Moreover, $v_{1},\dots,v_{\ell}$ is a unique path of length $\ell-1$ from the root $v_{1}$. Note that other edges in $T^{\prime}$ are the edges of the original tree $T$. In this way we come to the following conclusion.
@@
The intersection of an area of an extension path and the original graph is a tree of depth at most $\ell-1$.
Given a tree $T$ and its subtree $T^{\prime}$, there is a partial order on $V(T)\setminus V(T^{\prime})$ defined as follows: $u\leq^{T^{\prime}}v$ if $u$ belongs to the path from $v$ to $T^{\prime}$. For each $v\in V(T)\setminus V(T^{\prime})$ the set $\{u:v\leq^{T^{\prime}}u\}$ induces a tree ${\mathcal{I}}_{v}^{T^{\prime}}$ rooted at $v$. We will not mention $T^{\prime}$ in the exponent when it is clear from the context.
@@
Let $T_{0},T_{1},\dots,T_{k}$ be an $\ell$-growth sequence for an amoeba $A=(H,m)$ and $v\in V(T_{k})\setminus V(T_{0})$ such that ${\mathcal{I}}_{v}^{T_{0}}$ has depth at most $\ell-1$. If for some vertex $u$ in ${\mathcal{I}}_{v}^{T_{0}}$ there are two different paths $P$ and $P^{\prime}$ of equal lengths from $u$ to a leaf, then
$\text{{diam}}(H)\leq 2\ell-4\ \text{and}\ \text{{diam}}(T_{0})\leq 2\ell-6.$
@@
W.l.o.g. we can assume that $P$ and $P^{\prime}$ have no common edges. Indeed, otherwise we can change the vertex $u$ by the last common vertex of $P$ and $P^{\prime}$. Choose the maximal $i$ such that at least one edge of $P\cup P^{\prime}$ is not in $T_{i}$. W.l.o.g. we can assume that $e=(u_{0},u_{1})$ is the last edge of $P$ and $e$ is not in $T_{i}$. It means that $e$ is added in a growth in $T_{i}$ of a copy $A=(H,m)$ of the amoeba. So an extension of $A$ contains an extension path from a root $r$ of $A$ to $u_{1}$. We claim that $u$ is not on the path from $r$ to $u_{1}$ that has no common edges with $P^{\prime}$. Indeed, otherwise $P^{\prime}$ does not intersect $H$ and the extension path can be changed to the path containing $P^{\prime}$, contradicting the minimality condition; see Fig. 3(a). Also, $r$ is not on $P$. Indeed, otherwise the extension should add more edges, since the length of the whole path $P$ is less than $\ell$; see Fig. 3(b).
@@
  1. (i)
    vertices $u$, $u_{1}$, $r$ are end-points of paths starting in a vertex $w$; see Fig. 3(c);
  2. (ii)
    the path from $r$ to $u_{1}$ has a common edge with $P^{\prime}$; see Fig. 3(d).
In both cases $A$ is contained in the subtree rooted at $r$ since the path from $r$ to $u_{1}$ has no common edges with $A$. This subtree is inside ${\mathcal{I}}^{T_{0}}_{u}$ so its depth is at most $\ell-2$, which implies that diameter of $H$ is at most $42\ell-$ 4. Also, any path starting in $r$ and containing $w$ and $u$ should have length at most $\ell-1$, by the minimality condition. Since $T_{i}$ contains $T_{0}$ and the path from $r$ to $T_{0}$ has at least 3 edges, the diameter of $T_{0}$ is at most $62\ell-$ 6. ∎
Lemma 2 puts restrictions on extensions using ‘new’ edges only. Non-isomorphic $\ell$-growths are possible in generation sequences even if $H$ has the diameter at least $62\ell-$ 6; see an example in Fig. 4.
@@
Figure 5: A not mortal amoeba with one root with respect to $2$-extensions. The next step uses the left lower $P_{5}$ for extensions. Repetitions of the process continue forever
It is clear that the last tree in any generation sequence is confining. Thus, mortality implies the existence of a confining tree. The converse is proven only for $\ell\in\{1,2\}$ and it remains open in general.
@@
The ‘if part’ is obvious. The ‘only if part’: take a confining tree and a copy of the amoeba. Start a generation sequence from this copy. Note that $1$- and $2$-growths of a copy of the amoeba in any tree are isomorphic.
Indeed, if an extension path from a root contains a new edge then the root should be a pendant vertex in the initial tree or should be adjacent to a pendant vertex since $\ell\leq 2$. Thus, we add either a pendant edge or a path of length $2$. In the tree that is the union of two different growths these edges or paths belong to the same orbit of automorphisms. Therefore, the different growths are isomorphic to each other. This implies that at each step of extension new edges can be chosen from the confining tree. Therefore, the sequence is finite. ∎
@@
In the general case it is open whether the immortality is the negation of mortality. For the extension parameter $\ell\in\{1,2\}$ it is true due to Lemma 3. If an amoeba is not immortal then the last tree of its finite generation sequence is a confining tree for the amoeba. Thus, all its generation sequences are finite. So the amoeba is mortal.
In the proof of the next theorem we will use the notion of full $t$-ary tree of depth $\ell$. To this end we denote by $T^{t,\ell}$ a rooted tree in which every node at distance less than $\ell$ from the root has $t$ children, while every node at distance $\ell$ is a leaf.
@@
An amoeba is mortal w.r.t all trees if and only if it is totally confined, regardless of its extension parameter.
@@
In one direction it is easy. Suppose an amoeba is mortal w.r.t all trees. Take a tree $T$ and any $\ell$-growth sequence $T=T_{0},\dots,T_{k}$ starting in $T$. It is finite and the last tree $T_{k}$ in the sequence confines the amoeba and contains $T$.
For the other direction, take a tree $T_{0}$ and a totally confined amoeba $A=(H,m)$. Let $\delta=\max_{v\in V(H)}\big{(}d(v)+m(v)\big{)}$, $d=\mathop{\mathrm{diam}}(T_{0})$. Choose a confining tree $T$ that contains a full $\delta$-ary tree $T^{\delta,3\ell+2d}$ of sufficiently large depth, say, $3\ell+2d$. Choose a copy of $T_{0}$ containing the center of $T^{\delta,3\ell+2d}$. Start $\ell$-growth sequence $T_{0},T_{1},\dots$ from this particular copy. If $\mathop{\mathrm{diam}}(T_{i})\leq 2\ell-6$ then any copy of $A$ in it is at distance at most $62\ell-$ 6 from the center of $T^{\delta,3\ell+2d}$. Thus, any growth of $T_{i}$ can be embedded to $T$. If an $\ell$-growth of $T_{i}$ uses edges of $T_{0}$, then the corresponding copy of $A$ is at distance at most $2d$ from the center of $T^{\delta,3\ell+2d}$. Thus, again, the growth can be embedded to $T$. If an $\ell$-growth of $T_{i}$ does not use edges of $T_{0}$ and $\mathop{\mathrm{diam}}(T_{i})>2\ell-6$, then Lemma 2 implies that all such growths are isomorphic. So we apply the same arguments as in the proof of Lemma 3 and conclude that $T_{i+1}$ can be embedded in $T$. Therefore, the sequence is finite. ∎